The width of a rectangle is 18 in. less than its length. If the area of the rectangle is 1440 in^2, find the dimensions of the rectangle.
Answer by nyc_function(2741)

width = x - 18
length = x
area = 1440
You need the equation: A = L*W
1440 = x(x - 18)
Can you take it from here?
nycfunction@yahoo.com
Guido